# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v1

@@ 4.4 Attacking the Surrogate Classifier
In this section, we aim to answer RQ3, i.e., if we can jailbreak aligned LLMs by attacking the surrogate classifier. This section is symmetrical to Section 4.3. Instead of applying the attack on the LLMs and evaluating the inputs on the candidate classifiers, we attack the latter and evaluate how well the adversarial inputs transfer to the LLMs.
Baseline ASR on candidates. 9(a) reports the ASR of the attack on the candidate classifiers, which corresponds to the percentage of unsafe modified samples from the test dataset that were labeled safe by the candidate classifier (using the best threshold defined above). We note that when the candidate classifier uses the entire model, the corresponding ASR is close to the baseline ASR (e.g., 20% for Llama 2).
Transfer to the LLM. 9(b) shows the proportion of adversarial inputs crafted on the candidate classifiers that induce a misclassification on the model. Interestingly, the results do not exhibit a strongly increasing trend with respect to the candidate size like in the previous sections. For three of the models, we observe a peak above 70% when the normalized candidate size approaches 60%. In addition, attacking the candidate classifiers and transferring the resulting inputs to the model seem to work better than attacking the model directly. For example, using 50% of Llama 2, it is possible to achieve an ASR of 70%, well above the baseline ASR in Table 3.
@@
Takeaway 5: Adversarial objective of attacks White-box attacks have relied on heuristic-driven adversarial objectives (e.g., maximize the probability of a specific target output). By attacking the surrogate classifier, the adversarial objective is misclassification: a heuristic-free objective that removes unnecessary constraints on the search space of adversarial inputs.
Efficiency gains. Attacking the surrogate classifier has lower computational costs than attacking LLM, as it depends on the candidate size $\delta$ (of which the memory requirements scale linearly with). Paired with the previous results, attacking the surrogate classifier 9(b) shows that it is possible to achieve a transferability rate above 70% for nearly all models with a normalized candidate size of only 50%. The original attack requires a NVIDIA A100 with 40 GB for all models (at half-precision). Attacking 50% of the model (as presented in our work) means that under 20 GB of VRAM will be sufficient, enabling research on consumer-grade GPUs (e.g., NVIDIA RTX 4090). For industry-grade GPUs (e.g., NVIDIA A100, H100, etc.), this result enables in-depth study on hyperparameters of the attack such as the batch size.
Takeaway 6: Efficacy and efficiency of attacks Extracting the safety classifier of an LLM results in an increased attack efficacy (higher ASR) and efficiency (smaller model to attack) when attacking the surrogate classifier and transferring the modified inputs to the model. For example, attacking Llama 2 with only 50% of the model architecture leads to an ASR of 70%, well above the baseline ASR of 22%.
@@ 5.2 Implication for White-box Attacks
Classification objective. As opposed to evasion attacks on classifiers which cause misclassifications, jailbreak attacks aim to induce unsafe behaviors. This objective is more complex because it follows from the safety definition. When extracting the classifier, we explicitly uncover the decision made by the model. This gives a clearer signal which does not rely on heuristics (e.g., maximizing a target sequence that is likely to induce a jailbreak) and increases the attack efficacy.
Efficiency. Recent white-box gradient-based attacks on LLM are based on the GCG attack [61]. Although there have been significant performance improvements over time, these attacks compute gradients and, therefore, they are limited when the models are too large. Our results from Section 4.4 suggest that investigating the robustness of alignment through attacks is reducible to attacking the surrogate classifier. This implies fewer computations and thus higher efficiency (and scalability) of attacks. Our approach was applied on LLMs with less than 9 billion parameters and showed that using 50% of it is enough for the attack. If this result scales to larger models (ranging from 13 billion to 400 billion parameters), it would enable the application of white-box attacks on those models.
@@ 5.3 Safety Classifier
Start point of the classifier. We focus on the structures $f_{1,1+\delta}$, i.e., contiguous sets of decoders starting with the first layer. This allowed to locate the end of the safety classifier by evaluating the performance of the corresponding candidates. However, it may be that the classifier does not start at the first layer. Therefore, studying structures $f_{i,i+\delta}$ with $i>1$ may be of interest to refine the extraction but introduce a new dimension of complexity. Since the input space of decoder $i$ must match the output space of decoder $i-1$, new techniques need to be introduced to allow the creation of candidates that truly use the information of the structure. For example, a candidate classifier based on $f_{i,i+\delta}$ should not use information from $f_{1,i-1}$ or learn new information outside the mapping from structure representations to classification. We leave such explorations to future work.
Finer-grained extraction. Our approach uses results from previous work on the model representation of unsafe and safe inputs to reduce the space of structures. Within decoder blocks are multiple attention heads with different patterns learned. Therefore, it is possible to further refine the extraction of the classifier by removing attention heads that do not contribute to classification. The intractable number of possibilities implies the need for heuristics to search the space of structures efficiently. Since the candidates use a simple linear classification head, its performance is a function of how linearly separable the two classes are. Thus, a greedy approach in which attention heads are selected on the basis of how much they separate the embeddings could be a first step toward a finer extraction. However, not selecting certain attention heads of a decoder implies that the output space is different, which could lead to the issues mentioned above.
Multiclass. In this work, we considered the problem as a binary classification problem: classifying unsafe and safe inputs. As discussed in Section 2.2, multiple taxonomies of unsafe behaviors have been introduced. These taxonomies allow for better evaluations of attacks and give a finer-grained understanding of where alignment fails (e.g., aligned models might be more prone to accept prompt from the “false advertising” category [47]). Thus, a reasonable extension of this work would be to study the classification between categories. By studying this through a multiclass perspective, the task of the classifier changes from identifying inputs to be refused to identifying why these inputs should be refused. The challenge that emerges is to identify the predicted label from the LLMs as they tend to have generic refusal outputs, making the mapping from the output to an unsafe category hard, if not infeasible.
Labels. Part of our approach involves obtaining the classifications from the LLM (that correspond to the decision of the corresponding safety classifier). This involves two steps: generating the output from the LLM, then mapping this output text to a class, using the classification rule $\mathcal{R}$ defined in Section 3.3. LLMs are deployed in a way that enables nondeterministic outputs, leading to probabilities of refusal and compliance. To ensure reproducibility and reduce computational costs, we set the temperature to 0. However, studying the temperature and how it influences the probabilities of labels would lead to a more precise understanding of safety classifier and its decision boundary. We leave this study to future work.
@@ 6.1 Jailbreak and Representations
Improving the safety of LLMs is one of the most prevalent research areas: numerous publications have studied the link between jailbreak and model representations. It is necessary that alignment creates a linear separation between refused and accepted prompt for some representations, since the output is explicitly linearly dependent on the last embedding [24]. Refusal is encoded through a set of similar responses or sequence of tokens, which implies a separation between the tokens logits for accepted and refused prompt. Several papers have shown the distinction at the decoder level [24, 21, 56, 19, 53] or at the head of the attention level [57]. In addition, previous work has also shown that representations can be used to manipulate the model  [59] and increase its robustness at a negligible utility cost  [60]. Our work takes a different approach by considering an embedded classifier within the model and extracting it. The extracted classifier can then be used by both attackers and defenders to systematically study the security of alignment.
@@ 6.3 Pruning
In addition to studying their safety, an active area of research on LLMs is pruning: reducing the size of the model while maintaining most of its capabilities [10, 28]. For LLMs, pruning can be divided into two categories: width pruning and depth pruning. The former aims to reduce the size of the projection matrices, while the latter aims to remove decoders. The rise of security concerns led to the intersection of pruning and security. Recent work has shown that safety alignment is inherently low rank and can be isolated [45] and that pruning can offer higher safety [13]. Our work extracts the safety classifier within the model; therefore, the objective is different. Instead of reducing the size of the model while maintaining capabilities, which results in some loss of information, we aim to extract the safety classifier. Thus, minimizing the information lost on its decision boundary by having as little training as possible is crucial.
@@ 7 Conclusion
In this paper, we hypothesized that alignment embeds a safety classifier within LLMs, and we introduced techniques to extract best candidate classifiers for it. We empirically verified the agreement with the models of the best candidates in both benign and adversarial settings. We also demonstrated how the extracted classifier can be used to systematically study alignment security, as the candidate classifiers both substantially improved the efficiency (smaller models) and efficacy (higher success rate) for attackers and defenders. Finally, while we studied alignment in a safety context, hallucination, dishonesty, bias, and lack of fairness are all other examples of failures of alignment and our findings could also potentially benefit their understanding.
@@ Acknowledgments
The authors thank Chaowei Xiao and Rahul Chatterjee for their helpful comments in previous iterations of the work.
@@
Open Science
We comply with the open science policy by committing to share the following code and research artifacts of this paper:
  • The entirety of our code along with the corresponding set-up and dependencies environment for each part of the experiments. This includes the code to recreate the intermediate datasets, the training of the classification heads, and the code used to attack the model classifier estimations. This allows other researchers to reproduce our methodology and our results from scratch, but as this can be time and compute-intensive, we additionally release the following intermediary artifacts as well.
  • The intermediate datasets with the prompts, embedding representations, and labels for each model in benign and adversarial settings. This allows reproducing all results from Section 4.2, Section 3.1 and Section A.3 with limited computing resources.
  • The artifacts of the attack experiments. This includes in particular the perturbation and the corresponding classification by the model or the classifier estimations for each sample.
Ethics Considerations
This work studies the alignment of LLMs, specifically safety. Our primary objective is to demonstrate that alignment embeds a safety classifier into LLMs, responsible for deciding between refusal or compliance, and that extracting it can be beneficial from both an attack and defense perspective due to the improvements in both efficiency (smaller model to consider) and efficacy (higher attack success rate).
To be clear, in this paper, we offer a new paradigm to reason about LLM alignment in the context of jailbreak attacks. By now, this is a very well-known limitation that aligned LLMs are susceptible to these type of input perturbations. Our work does not raise any potential ethical concern related to the finding and disclosure of new jailbreak vulnerabilities, as we only study existing ones.
The potential harms that arise from similar type of studies to ours is (1) if it were evaluated against real deployed LLM services, and (2) if the paper were to contain information possibly unsafe to some. We describe next how we mitigate these non-existent concerns for our work.
First, recall from Section 3 that to demonstrate that alignment embeds a safety classifier, our methodology requires white-box access to the LLM. As a result, our approach is exclusively evaluated on open-source LLMs and at no point deployed against proprietary or API-gated online services (such as GPT, Claude, etc.).
Finally, our paper does not contain any unsafe information of itself whatsoever. We find it unnecessary for the purpose of our explanations to show in the body of our paper any specific examples of unsafe prompt or information from our datasets. However, some research artifacts mentioned in Open Science (e.g., intermediate datasets) contain adversarial examples that could lead to unsafe information generated by the corresponding LLM. We find this risk to be acceptable as it (a) allows for the verification of our results as per our pledge to open science, (b) can be communicated through the documentation released with our code, and (c) would be expected to be found by anyone looking into the artifacts for our paper (similarly to anyone using the established datasets that we work with).
@@ Appendix A Appendix
(b) Gemma 2
(c) Granite
(d) Llama 2
(e) Qwen 2.5
@@
(b) Gemma 2
(c) Granite
(d) Llama 2
(e) Qwen 2.5
@@ A.1 Confusion Matrices
Figure 10 and Figure 11 show the confusion matrices of the 5 models after applying the GCG attack on the LLM.
@@ A.2 Attacking Safe Inputs
Setup. We also augment the dataset by adding targets corresponding to a refusal. Different aligned models have different refusal answers depending on how they were trained. However, we saw limited impact on the choice of the target string on the performance of the attack and kept the same target for all prompts and models: “I cannot fulfill your request. I’m just an AI".
Safe ASR. Similarly to Section 4.3, the attack is applied to the models, this time with safe inputs (with the objective of making the model refuse). Table 4 reports the ASR for all settings. We first see that the highest ASR corresponds to Llama 2 (99%), which has the lowest ASR on unsafe inputs (12% to 22%). In contrast, attacking Qwen 2.5 produced a high ASR for unsafe and safe inputs. Finally, Gemma 1 and Gemma 2 have shown medium ASR on unsafe inputs, also exhibited for safe inputs.
@@
Table 4: Attack success rates (ASR) of the models after applying the attack on safe inputs.
Results. 12(a) and 12(b) report the ASR on candidate classifiers and the transferability rate to the LLM, for safe inputs. We notice the same trends on the transferability rate with a peak for most models with an exception for Llama 2. Furthermore, for OR-Bench, the ASR on the candidate classifiers corresponds exactly to the transferability rate, implying that the success of the attack is entirely determined by how well it performs on the candidate classifier. This finding aligns with the performance in adversarial setting in 8(b): Llama 2 candidate classifiers have the most stable performance on adversarial examples from the LLM compared to other models.
@@ A.3 Other Models
Table 5: $F_{1}$ score of the other models on the two datasets.
Figure 13: Silhouette score for other models.
(a) Llama 3
(b) Mistral
(c) Zephyr RMU
Figure 14: Confusion matrices of other models on AdvBench
(a) Llama 3
(b) Mistral
(c) Zephyr RMU
Figure 15: Confusion matrices of other models on OR-Bench
Results. Table 5 reports their accuracy and F1 score on the two datasets. All models achieve an F1 score lower than 0.8 on OR-Bench, translating to a weaker performance according to the setup. Figure 16 and Figure 17 show the performance of the candidates in benign settings, with the same setup as Section 4.2. We see that the trend is similar in 16(a) with a convergence after a certain candidate size. However, the F1 score is lower than the other models studied. Further, 16(b) shows a different trend, with a decrease at the middle of the model. Figure 17 shows the cross-dataset experiment for the other models studied. Figure 14 and Figure 15 show the confusion matrices in benign settings for the other models studied.
@@
Possible causes. The Mistral model is described as a ”quick demonstration" that ”does not have any moderation mechanisms" [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 in the ”Limitations” section [17]. Llama 3, similarly to Llama 2, did go through alignment. However, the tone of the output was adjusted, specifically for refusal outputs. Therefore, detecting the classification made by the model is more challenging. Zephyr RMU is a model in which unsafe information has been removed through unlearning [20], therefore the lower performance is due to the way labels are assigned: in our experimental setup, we assign labels based on whether the model complies with the input prompt or not. Thus, since Zephyr RMU underwent unlearning, it may still comply with unsafe input prompts with a safe output. We argue that models that are made more secure through unlearning are orthogonal to the scope of the paper: this method likely removes part of the safety classifier alongside the unsafe knowledge, preventing its extraction.

### v2

@@ 4.4 Attacking the Surrogate Classifier
In this section, we aim to answer RQ3, i.e., if we can jailbreak aligned LLMs by attacking the surrogate classifier. In this section, we attack the candidate classifier and evaluate how well the adversarial inputs transfer to the LLMs.
Figure 9: Transferability rate to the LLM after applying the white-box attack on the candidate classifiers.
Transfer to the LLM. Figure 9 shows the proportion of adversarial inputs crafted on the candidate classifiers that induce a misclassification on the model. Interestingly, the results do not exhibit a strongly increasing trend with respect to the candidate size like in the previous sections. For three of the models, we observe a peak above 70% when the normalized candidate size approaches 60%. In addition, attacking the candidate classifiers and transferring the resulting inputs to the model seem to work better than attacking the model directly. For example, using 50% of Llama 2, it is possible to achieve an ASR of 70%, well above the baseline ASR in Table III.
@@
Takeaway 5: Adversarial objective of attacks White-box attacks have relied on heuristic-driven adversarial objectives. Using misclassification as the objective removes unnecessary constraints on the search space of adversarial inputs.
Figure 10: Runtime and VRAM usage to craft one adversarial example depending on the candidate size. The dashed line represents the averaged baseline runtime (840s) and VRAM usage (31 GB) when attacking the LLM directly.
Efficiency. Figure 10 shows the runtime [^4] This does not include the time to train the candidate classifier as it is negligible, taking less than a minute and happening once for all samples. and VRAM usage (i.e., peak GPU memory occupied) when applying the attack on one sample, averaged over 100 samples. We see that attacking the surrogate classifier has lower computational costs than attacking LLM, as it depends on the candidate size $\delta$ (which scales linearly with the VRAM usage like the runtime, as seen in Figure 10). We note that both the VRAM usage and the runtime do not not reach the baseline averages of 31 GB and 840s. We attribute this to the adversarial objective of the baseline GCG attack: it requires the attack to generate an output sequence at each step to compare to the target sequence, directly impacting its efficiency.
Paired with the previous results from Figure 9 in which all models achieve a transferability rate above 70% with only half of the LLM architecture, this highlights the efficacy and efficiency of our approach. This also shows that studying the robustness of LLMs can be made scalable.
Indeed, the original attack requires a NVIDIA A100 with 40 GB for all models (at half-precision). Attacking 50% of the model (as presented in our work) means that 20 GB of VRAM will be sufficient, enabling research on consumer-grade GPUs (e.g., NVIDIA RTX 4090). For industry-grade GPUs (e.g., NVIDIA A100, H100, etc.), this result enables in-depth study on hyperparameters of the attack.
Takeaway 6: Efficacy and efficiency of attacks Attacking the surrogate classifier leads to a significant improvement in efficacy (ASR) and efficiency (runtime and VRAM usage). For example, attacking Llama 2 with only 50% of the LLM leads to an ASR of 70%—well above the baseline ASR of 22%— with half the runtime and VRAM usage.
@@ 5.2 Implication for White-box Attacks
Classification objective. As opposed to evasion attacks on classifiers which cause misclassifications, jailbreak attacks aim to induce unsafe behaviors. This objective is more complex because it follows from the definition of safety. When extracting the classifier, we explicitly uncover the decision made by the model. This gives a more accurate signal that does not rely on heuristics (e.g., maximizing a target sequence that is likely to induce a jailbreak) and thus leads to a higher attack efficacy. We also note that such heuristics on the adversarial objective not only implicitly reduce the efficacy, can also reduce the efficiency (see Section 4.4) as it requires multiple forward pass to the entire model to compute an output sequence.
Efficiency. Recent white-box gradient-based attacks on LLM are based on the GCG attack [65]. Although there have been significant performance improvements over time, these attacks compute gradients and, therefore, they are limited when the models are too large. Our results from Section 4.4 suggest that investigating the robustness of alignment through attacks is reducible to attacking the surrogate classifier. This implies fewer computations, and thus higher efficiency (and scalability) of attacks. Our approach was applied on LLMs with less than 9 billion parameters and showed that using 50% of them is enough for the attack. A similar efficiency could be obtained from a transfer attack using a smaller LLM as the source. However, this would come with a lower efficacy and the results might be less informative on the failure points of the target model since a different model is used.
Attack scenarios. Our work can be applied to any jailbreak attack (albeit more suited for optimization-based attacks), making it extensible to multiple scenarios. For instance, in the presence of a guard model that prevents unsafe outputs (e.g., LlamaGuard [18]), recent work has shown that an attack can be “propagated” to this guard model [28]. Our work can be extended to such scenarios by extracting the safety classifiers of (a) the LLM and (b) the guard model, therein improving efficiency.
Black-box. We limit the threat model to white box, since we need access to the LLM’s weights to obtain the surrogate classifier. However, a promising direction would be to study the transferability of safety classifier of a source LLM onto that of a target LLM. Transferability of white-box attacks is a well-known phenomenon of adversarial machine learning [11] also present for jailbreak attacks [65, 27]. An analysis with candidate classifiers could give a finer-grained understanding of the transferability of jailbreak attacks.
Defenses. Two axes are available to defend against our approach. First, as any attack against an ML model, adversarial training [51] remains the de facto standard for increasing ML robustness. Instead of applying such training to the entire model (which would be highly inefficient given its size), it can be applied to the safety classifier with the classification objective. For a finer-grained threat model where the adversary is computationally bounded (by either runtime or memory), explicitly training the LLM to increase the safety classifier size (e.g., by modifying the learning objective) would limit the feasibility of the attack due to computational constraints. This, however, incurs a trade-off for other defenses such as detection of adversarial examples in the intermediate layers [49] as the detection would happen at the end of the safety classifier.
@@ 5.3 Safety Classifier
Start point of the classifier. We focus on the structures $f_{1,1+\delta}$, i.e., contiguous sets of decoders starting with the first layer. This allowed to locate the end of the safety classifier by evaluating the performance of the corresponding candidates. However, it may be that the classifier does not start at the first layer. Therefore, studying structures $f_{i,i+\delta}$ with $i>1$ may be of interest to refine the extraction but introduce a new dimension of complexity. Since the input space of decoder $i$ must match the output space of decoder i1i-1, new techniques need to be introduced to allow the creation of candidates that truly use the information of the structure. For example, a candidate classifier based on $f_{i,i+\delta}$ should not use information from $f_{1,i-1}$ or learn new information outside the mapping from structure representations to classification. We leave such explorations to future work.
Labels. Part of our approach involves obtaining the classifications from the LLM. This involves two steps: generating the output from the LLM, then mapping this output text to a class, using the classification rule $\mathcal{R}$ defined in Section 3.3. LLMs are deployed in a way that allows nondeterministic outputs, leading to probabilities of refusal and compliance. To ensure reproducibility and reduce computational costs, we set the temperature to 0. However, studying the temperature and how it influences the probabilities of labels would lead to a more precise understanding of safety classifier and its decision boundary. We leave this study to future work.
Finding optimal depth. From an adversarial perspective, finding the position of the safety classifier is crucial, as it can be expensive to run the attack on various candidates. A first strategy to pinpoint the most likely optimal depth would be to follow the procedure of Section 3.1, i.e., using the silhouette score. However, the adversary may be constrained in resources and have limited GPU memory. In that case, multiple strategies are possible, such as starting at the maximum candidate size for the GPU memory and lowering it or doing a binary search on the candidate size.
@@ 6.1 Jailbreak and Representations
Improving the safety of LLMs is one of the most prevalent research areas: numerous publications have studied the link between jailbreak and model representations. It is necessary that alignment creates a linear separation between refused and accepted prompt for some representations, since the output is explicitly linearly dependent on the last embedding [26]. Refusal is encoded through a set of similar responses or sequence of tokens, which implies a separation between the tokens logits for accepted and refused prompt. Several papers have shown the distinction at the decoder level [26, 23, 59, 21, 56] or at the head of the attention level [60]. In addition, previous work has also shown that representations can be used to manipulate the model  [63] and increase its robustness at a negligible utility cost  [64]. These works often use linear probing to obtain “probes” that classify the representations of prompts. However, the utility of these probes remains limited: they cannot be used to produce adversarial examples as they do not hold any gradient information from embeddings to tokens (i.e., from intermediary output to input). Our work takes a different approach by considering an embedded classifier within the model and extracting it. We leverage linear probing to map the output of a structure (a subset of the LLM) to a classification. This allows the application of jailbreak attacks to produce token-level adversarial examples (i.e., unsafe input prompts that cause the LLM to comply). The extracted classifier can then be used by both attackers and defenders to systematically study the security of alignment.
@@ 6.3 Pruning
In addition to studying their safety, an active area of research on LLMs is pruning, i.e., reducing the size of the model while maintaining most of its capabilities [13]. For LLMs, pruning can be divided into three axes: depth, width, and length pruning. Depth pruning focuses on removing layers of the model, width pruning aims to reduce the size of the projection matrices, and length pruning focuses on cutting tokens of the input stream and avoiding redundant computations. As those axes are orthogonal, they can be combined to obtain better results [43]. The rise of security concerns led to the intersection of pruning and security. Recent work has shown that safety alignment is inherently low rank and can be isolated [48] and that pruning can offer higher safety [15]. Although these techniques could achieve a similar objective, they generally involve modifying weights from the selected parts of the model. As we aim to extract the safety classifier, minimizing the information lost on its decision boundary by having as little training as possible is crucial, which is why we chose the least invasive method (linear probing). This is also the most efficient way to extract the classifier, as only the necessary weights are trained, i.e., those that map the representation to the label.
@@ 7 Conclusion
In this paper, we hypothesized that alignment embeds a safety classifier within LLMs, and we introduced techniques to extract best candidate classifiers for it. We empirically verified the agreement with the models of the best candidates in both benign and adversarial settings. We also demonstrated how the extracted classifier can be used to systematically study alignment security, as the candidate classifiers substantially improved both the efficiency (lower runtime and VRAM usage) and the efficacy (higher success rate) for attackers and defenders. Finally, while we studied alignment in a safety context, other failures of alignment such as hallucination, dishonesty, bias, and lack of fairness could also benefit from our approach.
@@ Appendix A Appendix
(b) Granite
(c) Llama 2
(d) Qwen 2.5
@@
(b) Granite
(c) Llama 2
(d) Qwen 2.5
@@ A.1 Confusion Matrices
Figure 11 and Figure 12 show the confusion matrices of the 4 models after applying the GCG attack on the LLM.
@@ A.2 Attacking Safe Inputs
Setup. We also augment the dataset by adding targets corresponding to a refusal. Different aligned models have different refusal answers depending on how they were trained. However, we saw limited impact on the choice of the target string on the performance of the attack and kept the same target for all prompts and models: “I cannot fulfill your request. I’m just an AI”.
Safe ASR. Similarly to Section 4.3, the attack is applied to the models, this time with safe inputs (with the objective of making the model refuse). Table IV reports the ASR for all settings. We first see that the highest ASR corresponds to Llama 2 (99%), which has the lowest ASR on unsafe inputs (12% to 22%). In contrast, attacking Qwen 2.5 produced a high ASR for unsafe and safe inputs. Finally, Gemma 1 has shown medium ASR on unsafe inputs, also exhibited for safe inputs.
@@
TABLE IV: Attack success rates (ASR) of the models after applying the attack on safe inputs.
Results. 13(a) and 13(b) report the ASR on candidate classifiers and the transferability rate to the LLM, for safe inputs. We notice the same trends on the transferability rate with a peak for most models with an exception for Llama 2. Furthermore, for OR-Bench, the ASR on the candidate classifiers corresponds exactly to the transferability rate, implying that the success of the attack is entirely determined by how well it performs on the candidate classifier. This finding aligns with the performance in adversarial setting in 8(b): Llama 2 candidate classifiers have the most stable performance on adversarial examples from the LLM compared to other models.
@@ A.3 Other Models
TABLE V: $F_{1}$ score of the other models on the two datasets.
Results. Table V reports their accuracy and F1 score on the two datasets. All models achieve an F1 score lower than 0.8 on OR-Bench, translating to a weaker performance according to the setup. Figure 14 and Figure 15 show the performance of the candidates in benign settings, with the same setup as Section 4.2. We see that the trend is similar in 14(a) with a convergence after a certain candidate size. However, the F1 score is lower than the other models studied. Further, 14(b) shows a different trend, with a decrease at the middle of the model. Figure 15 shows the cross-dataset experiment for the other models studied.
@@
Possible causes. The Mistral model is described as a ”quick demonstration” that ”does not have any moderation mechanisms” [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 in the ”Limitations” section [19]. Llama 3, similarly to Llama 2, did go through alignment. However, the tone of the output was adjusted, specifically for refusal outputs. Therefore, detecting the classification made by the model is more challenging. Zephyr RMU is a model in which unsafe information has been removed through unlearning [22], therefore the lower performance is due to the way labels are assigned: in our experimental setup, we assign labels based on whether the model complies with the input prompt or not. Thus, since Zephyr RMU underwent unlearning, it may still comply with unsafe input prompts with a safe output. We argue that models that are made more secure through unlearning are orthogonal to the scope of the paper: this method likely removes part of the safety classifier alongside the unsafe knowledge, preventing its extraction.